Please solve the system of the linear equation using any algebraic method and if possible show a solution check.
3y=11-2x
3x=y-11
Thank you

Answer by addingup(3677)

Substitution:
3y = 11-2x; 3y+2x = 11
3x = y-11; y = 3x+11 substitute for y in the equation above
:
3(3x+11)+2x = 11
9x+33+2x = 11
11x = -22
x = -2
:
3(-2)= y-11
-6+11 = y
y = 5
--------------------------------
Let's check our answers, substitute in the original equations:
x = -2
y = 5
3y = 11-2x
3(5) = 11-2(-2)
15 = 11+4 Correct
Next:
3x = y-11
3(-2) = 5-11
-6 = -6 Also correct.